# MC: Multi-concept Guidance for Customized Multi-concept Generation

### v2

@@ Abstract
Customized text-to-image generation aims to synthesize instantiations of user-specified concepts and has achieved unprecedented progress in handling individual concept. However, when extending to multiple customized concepts, existing methods exhibit limitations in terms of flexibility and fidelity, only accommodating the combination of limited types of models and potentially resulting in a mix of characteristics from different concepts. In this paper, we introduce the Multi-Concept guidance for Multi-Concept customization, termed MC, for improved flexibility and fidelity. MC decouples the requirements for model architecture via inference time optimization, allowing the integration of various heterogeneous single-concept customized models. It adaptively refines the attention weights between visual and textual tokens, directing image regions to focus on their associated words while diminishing the impact of irrelevant ones. Extensive experiments demonstrate that MC even surpasses previous methods that require additional training in terms of consistency with input prompt and reference images. Moreover, MC can be extended to elevate the compositional capabilities of text-to-image generation, yielding appealing results. Code will be publicly available at https://github.com/JIANGJiaXiu/MC-2.
Keywords:
Text-to-image generation Customized multi-concept generation Compositional generation1 Introduction
@@
Existing works performs customized multi-concept generation through joint training [25, 16] or merging single-concept customized models [47, 38, 29, 15]. Joint training [25, 16] requires access to the reference images of the specified concepts, and retrain all parameters when introducing a new concept. Merging single-concept customized models [47, 38, 29, 15] promisingly reduces the training cost. Nonetheless, it is challenging for the above methods to integrate multiple heterogeneous single-concept customized models, as they require specific network architectures and access to training data. Users may obtain a large amount of single-concept customized models from the community, but cannot acquire their training data in most cases. Therefore, these models are difficult to be integrated to generate compositions of multiple customized concepts. Albeit Cones-2 [30] does not require joint training or model merging for generating composition of customized concepts, it introduces additional and inconvenient condition signals, e.g., bounding boxes or segmentation masks. Encoder-based customization methods [32, 50] usually involve finetuning the pretrained text-to-image diffusion model to accept identity information extracted from the reference images, which are orthogonal to our work.
To address the above mentioned problems, we propose the Multi-Concept guidance for customized Multi-Concept generation, termed MC. Our proposed MC facilitates the seamless integration of separately trained heterogeneous single-concept customized models, allowing for the natural synthesis of a composition of their distinct concepts without additional training. As shown in Fig. 1, a Textual Inversion [13] model, a LoRA [21] and a DreamBooth [43] model are combined with our method, demonstrating its versatility without requiring extra training or additional conditioning information. Multi-concept guidance (MCG) acts as an avenue for the separately trained customized models to communicate with each other. In each step of MC, each customized model denoises the same latent noise map respectively, then MCG attempts to identify the activated regions of each concept and spatially disentangle them. When the activated regions of each concept have little overlap, the multiple customized concepts are more likely to be generated simultaneously with less incorrect attribute binding. Inspired by [46, 7, 35], we extract the cross-attention maps from the diffusion process. Certain cross-attention maps indicate the activated regions of each customized concept. Then MCG adaptively refines the attention weights between visual and textual tokens, directing image regions to focus on their associated words while diminishing the impact of irrelevant ones. Moreover, MC can be extended to enhance the compositional generation ability of existing text-to-image diffusion models, yielding appealing results.
@@
We propose a novel method MC for integrating various single-concept customized models to synthesize composition of different customized concepts, which does not require joint training, model merging or additional condition information.
With minor adjustments, MC can be extended to enhance the compositional generation ability of existing text-to-image diffusion models.
Extensive qualitative and quantitative evaluations demonstrate that our proposed MC significantly improves the performance of customized multi-concept generation and compositional text-to-image generation, even surpassing previous methods that require additional training.
@@ 2 Related Work
Customized multi-concept generation. Customized generation methods [13, 43, 49] aim to generate images of a specified concept given a few user-provided images of the concept. Custom Diffusion [25] first proposes to extend the task to a multi-concept scenario, where the compositions of multiple customized concepts are expected. Custom Diffusion finetunes a diffusion model given reference images of multiple concepts or merges single-concept customized diffusion models with constrained optimization. SVDiff [16] proposes the data augmentation strategy Cut-Mix-Unmix which splices two reference images together for finetuning the diffusion model. Some works [47, 38, 29] separately train customized models for each concept then merge the models to get one customized model with their proposed formulas. Cones [29] additionally requires finetuning for the merged customized model for better generation quality. Mix-of-Show [15] trains ED-LoRAs for each concept then merges them into one ED-LoRA via gradient fusion. Cones 2 [30] finetunes the text encoder for each concept, then derives a token embedding for it. Cones 2 only uses the embeddings for each concept at inference. Compositional Inversion [53] finetunes token embeddings for each concept then composes them via their proposed spatial inversion. Gen4Gen [51] proposes a data generation pipeline for training multi-concept customized models. Encoder-based customization methods FastComposer [50] and Subject-Diffusion [32] also support multi-concept customization, which are orthogonal to our work. VideoDreamer [8] and CustomVideo [48] consider multi-concept customization for text-to-video generation with diffusion models. DreamMatcher [34] enhances existing customization methods with semantic matching. Concurrent work [55] proposes to compose multiple LoRAs for image generation, but not particularly for multi-concept customization. Another line of work [3, 22, 2] focuses on extracting multiple concepts or attributes from a single image. Break-A-Scene [3] proposes a masked diffusion loss and a cross-attention loss to encourage the disentanglement between the concepts during training.
MC supports combining multiple separately trained customized models of different architectures, e.g. Textual Inversion [13], LoRA [21] and DreamBooth [43], without additional training or layout information.
@@
Sampling guidance for diffusion models. Guidance methods interfere with the sampling process of the diffusion models to achieve various effects. Some works [10, 19, 36, 20] improve the sampling quality of diffusion models or make the generation results more consistent with the input condition. Others [4, 24, 14, 9, 11] enable additional condition input for pretrained text-to-image diffusion models, to control layout, style or other attributes of the generated images.
@@ 3.2 Preliminaries
Equation 3 can act as the structure in Eq. 1, but it may confuse objects’ attributes or create an object that merges the specified concepts, as shown in the second row of Fig. 3 and the first row of Fig. 5. Composable LoRA [37] proposes to use the structure with LoRAs [44], i.e. to merge the concepts contained in different LoRA modules. As shown in Fig. 2, we adopt the architecture of multiple parallel diffusion models with different customized modules to start with. Limited by [28], the diffusion models are the instances of the same model. With this architecture, the customized modules can be of different forms, e.g. Textual Inversion [13], LoRA [21] and DreamBooth [43], as they are not required to be merged to one model, as shown in Fig. 1.
@@ 3.3 Multi-concept Guidance
Our proposed MC performs Multi-concept Guidance (MCG) at each step of the diffusion process. At each step of the diffusion process we refine the attention weights between visual and textual tokens in order to spatially disentangle the customized concepts. We define a loss objective to measure the extent to which the customized concepts are spatially disentangled. As illustrated in Fig. 2, the noised latent map is updated with the objective in the first stage. In the second stage, the diffusion models take the updated noised latent map as input and is produced via semantic merging for the next denoising step.
We attempt to measure the extent to which the customized concepts are spatially disentangled with the help of the cross-attention maps generated at each step of the diffusion process. They can be regarded as features showing how strongly each latent image patch is associated with each text token embedding. Each cross-attention layer in the unet takes latent image features and text features as input. After linear transformation , , , the attention maps are then:
@@
To spatially disentangle different customized concepts, we define another loss term to encourage less overlap of the regions activated by different customized modules. For sub-prompt , we first aggregate the attention maps of the trigger tokens to get one attention map. Here, we empirically average the attention maps. After being smoothed by a Gaussian filter, we get the final attention map indicating the activated region of the customized concept. Figure 3 visualizes the attention maps of the involved concepts. Our proposed inter-prompt disentanglement loss is then:
[…]
@@
[…]
@@ 4.1 Experiment Setup
Datasets. For customized multi-concept generation, we perform experiments on the CustomConcept101 dataset [25], consisting of 101 concepts and 101 unique compositions among the concepts. Each concept possesses 3 to 15 images. Each composition involves two concepts, accompanied by 12 prompts, e.g. photo of the {0} and {1}.
@@
Evaluation metrics. For customized multi-concept generation, subject fidelity and prompt fidelity are evaluated, following [43, 25]. For subject fidelity, we compute two metrics: CLIP-I [13] and DINO [43]. CLIP-I is the average cosine similarity between CLIP [39] embeddings of generated and real images. The DINO metric is the average pairwise cosine similarity between the ViT-S/16 DINO [6] embeddings of generated and real images. For a composition of two concepts, CLIP-I and DINO metrics are computed for each concept respectively then averaged. Prompt fidelity is measured as the average cosine similarity between prompt and image CLIP embedding, denoted as CLIP-T[43]. The placeholder token in the text prompt template is replaced with the proper concept category for extracting CLIP text feature.
@@
Baselines. For customized multi-concept generation, we compare our method with Custom Diffusion [25], Mix-of-Show [15] and Cones 2 [30]. Custom Diffusion optimizes the cross-attention key and value parameter in the unet plus a token embedding given reference images. It either jointly train for multiple concepts or combine multiple finetuned models into one via closed-form constrained optimization, denoted as Joint and Optimization respectively. Mix-of-Show trains ED-LoRA for each concept, then fuses multiple ED-LoRAs into one via gradient fusion. ED-LoRA contains two token embeddings and LoRA [21] for the diffusion model. Mix-of-Show introduces regionally controllable sampling to control the location of each concept, which requires user-input bounding boxes for each concept, thus is omitted in our experiment. Cones 2 finetunes the text encoder for each concept. Then a residual token embedding is derived from the text encoder for inference. Cones 2 proposes a layout guidance which requires user-input location information. The layout guidance is omitted in our experiment. Textual Inversion [13] and DreamBooth [43] is not included here as they are not intended for customized multi-concept generation. Their performance can be found in [25].
@@
For compositional generation, we compare our method with Attend-and-Excite [7], Divide-and-Bind [27] and Composable Diffusion [28]. Attend-and-Excite proposes a loss objective to maximize the attention values for each subject token. Divide-and-Bind proposes a attendance loss and a binding loss to better incorporate the semantics. For performance of Stable Diffusion v1-5 [42], please refer to [7].
@@
Implementation details. We use the Stable Diffusion v1-5 in our experiments. Without loss of generality, we choose LoRA as the single-concept customized model, as different single-concept customization methods basically finetune some selected part of the diffusion model with similar loss functions. We adopt a popular community implementation of LoRA 111https://github.com/kohya-ss/sd-scripts. LoRA modules are trained for linear layers and conv layers in text encoder and unet with rank set to 16. We do not finetune the token embeddings. All reference images are captioned as photo of a <concept name>. Each concept has a unique <concept name> defined by CustomConcept101, e.g. pet_dog1. All LoRAs are trained for 1000 steps with batch size set to 2, learning rate set to 1e-4. LoRA scale is set to 0.7 for merging the LoRA parameters into the diffusion model during inference. Unless mentioned otherwise, we use DPM++ sampler [31] with 30 sampling steps. Given a prompt template, e.g. photo of the {0} and {1}, we construct the sub-prompts for a composition as follows: (i) photo of the <cate1> and <cate2>, (ii) photo of the <concept1> and <cate2>, a <concept1>, (iii) photo of the <cate1> and <concept2>, a <concept2>. <cate1> is the category name of <concept1> defined by CustomConcept101, e.g. dog for pet_dog1.
| Method | CLIP-T | CLIP-I | DINO |
| Custom Diffusion (Joint) [25] | 0.789 | 0.645 | 0.339 |
| Custom Diffusion (Optimization) [25] | 0.788 | 0.646 | 0.339 |
| Mix-of-Show [15] | 0.728 | 0.670 | 0.394 |
| Cones 2 [30] | 0.827 | 0.624 | 0.264 |
| Ours | 0.767 | 0.686 | 0.406 |
| Method | Full Prompt | Min. Object | Text-Text |

### v3

# MC2: Multi-concept Guidance for Customized Multi-concept Generation

@@ Abstract
Customized text-to-image generation, which synthesizes images based on user-specified concepts, has made significant progress in handling individual concepts. However, when extended to multiple concepts, existing methods often struggle with properly integrating different models and avoiding the unintended blending of characteristics from distinct concepts. In this paper, we propose MC2, a novel approach for multi-concept customization that enhances flexibility and fidelity through inference-time optimization. MC2 enables the integration of multiple single-concept models with heterogeneous architectures. By adaptively refining attention weights between visual and textual tokens, our method ensures that image regions accurately correspond to their associated concepts while minimizing interference between concepts. Extensive experiments demonstrate that MC2 outperforms training-based methods in terms of prompt-reference alignment. Furthermore, MC2 can be seamlessly applied to text-to-image generation, providing robust compositional capabilities. To facilitate the evaluation of multi-concept customization, we also introduce a new benchmark, MC++. The code will be publicly available at https://github.com/JIANGJiaXiu/MC-2.
1 Introduction
@@
Existing works performs customized multi-concept generation through joint training [26, 17] or merging single-concept customized models [50, 39, 30, 16]. Joint training [26, 17] requires access to the reference images of the specified concepts, and retrain all parameters when introducing a new concept. Merging single-concept customized models [50, 39, 30, 16] promisingly reduces the training cost. Nonetheless, it is challenging for the above methods to integrate multiple heterogeneous single-concept customized models, as they require specific network architectures and access to training data. Users may obtain a large amount of single-concept customized models from the community, but cannot acquire their training data in most cases. Therefore, these models are difficult to be integrated to generate compositions of multiple customized concepts. Albeit Cones 2 [31] does not require joint training or model merging for generating composition of customized concepts, it introduces additional and inconvenient condition signals, e.g., bounding boxes or segmentation masks. Encoder-based customization methods [33, 54] usually involve finetuning the pretrained text-to-image diffusion model to accept identity information extracted from the reference images, which are orthogonal to our work.
To address the above mentioned problems, we propose the Multi-Concept guidance for customized Multi-Concept generation, termed MC2. Our proposed MC2 facilitates the seamless integration of separately trained heterogeneous single-concept customized models, allowing for the natural synthesis of a composition of their distinct concepts without additional training. As shown in Fig. 1, a Textual Inversion [14] model, a LoRA [21] and a DreamBooth [45] model are combined with our method, demonstrating its versatility without requiring extra training or additional conditioning information. Multi-concept guidance (MCG) acts as an avenue for the separately trained customized models to communicate with each other. In each step of MC2, each customized model denoises the same latent noise map respectively, then MCG attempts to identify the activated regions of each concept and spatially disentangle them. When the activated regions of each concept have little overlap, the multiple customized concepts are more likely to be generated simultaneously with less incorrect attribute binding. Inspired by [49, 8, 36], we extract the cross-attention maps from the diffusion process. Certain cross-attention maps indicate the activated regions of each customized concept. Then MCG adaptively refines the attention weights between visual and textual tokens, directing image regions to focus on their associated words while diminishing the impact of irrelevant ones. Moreover, MC2 can be extended to enhance the compositional generation ability of existing text-to-image diffusion models, yielding appealing results.
@@
We propose a novel method MC2 for integrating various single-concept customized models to synthesize composition of different customized concepts, which does not require joint training, model merging or additional condition information.
We construct a benchmark MC++ for evaluating multi-concept customization. The benchmark contains compositions of two to four concepts.
With minor adjustments, our method can be extended to enhance the compositional generation ability of existing text-to-image diffusion models.
Extensive qualitative and quantitative evaluations demonstrate that our proposed method significantly improves the performance of customized multi-concept generation and compositional text-to-image generation, even surpassing previous methods that require additional training.
@@ 2 Related Work
Customized multi-concept generation. Customized generation methods [14, 45, 53] aim to generate images of a specified concept given a few user-provided images of the concept. Custom Diffusion [26] first proposes to extend the task to a multi-concept scenario, where the compositions of multiple customized concepts are expected. Custom Diffusion finetunes a diffusion model given reference images of multiple concepts or merges single-concept customized diffusion models with constrained optimization. SVDiff [17] proposes the data augmentation strategy Cut-Mix-Unmix which splices two reference images together for finetuning the diffusion model. Some works [50, 39, 30] separately train customized models for each concept then merge the models to get one customized model with their proposed formulas. Cones [30] additionally requires finetuning for the merged customized model for better generation quality. Mix-of-Show [16] trains ED-LoRAs for each concept then merges them into one ED-LoRA via gradient fusion. Cones 2 [31] finetunes the text encoder for each concept, then derives a token embedding for it. Cones 2 only uses the embeddings for each concept at inference. Compositional Inversion [55] finetunes token embeddings for each concept then composes them via their proposed spatial inversion. Encoder-based customization methods FastComposer [54] and Subject-Diffusion [33] also support multi-concept customization, which are orthogonal to our work. VideoDreamer [9] and CustomVideo [51] consider multi-concept customization for text-to-video generation with diffusion models. DreamMatcher [35] enhances existing customization methods with semantic matching. Concurrent work [58] proposes to compose multiple LoRAs for image generation, but not particularly for multi-concept customization. Another line of work [3, 22, 2] focuses on extracting multiple concepts or attributes from a single image. Break-A-Scene [3] proposes a masked diffusion loss and a cross-attention loss to encourage the disentanglement between the concepts during training.
MC2 supports combining multiple separately trained customized models of different architectures, e.g. Textual Inversion [14], LoRA [21] and DreamBooth [45], without additional training or layout information.
@@
Sampling guidance for diffusion models. Guidance methods interfere with the sampling process of the diffusion models to achieve various effects. Some works [11, 18, 37, 20] improve the sampling quality of pretrained diffusion models or make the generation results more consistent with the input condition. Others [4, 24, 15, 10, 12] enable additional condition input for pretrained text-to-image diffusion models, to control layout, style or other attributes of the generated images.
@@ 3.2 Preliminaries
Equation 3 can act as the structure in Eq. 1, but it may confuse objects’ attributes or create an object that merges the specified concepts, as shown in the second row of Fig. 3 and the first row of Fig. 6. Composable LoRA [38] proposes to use the structure with LoRAs [46], i.e. to merge the concepts contained in different LoRA modules. As shown in Fig. 2, we adopt the architecture of multiple parallel diffusion models with different customized modules to start with. Limited by [29], the diffusion models are the instances of the same model. With this architecture, the customized modules can be of different forms, e.g. Textual Inversion [14], LoRA [21] and DreamBooth [45], as they are not required to be merged to one model, shown in Fig. 1.
@@ 3.3 Multi-concept Guidance
Our proposed MC2 performs Multi-concept Guidance (MCG) at each step of the diffusion process. At each step of the diffusion process we refine the attention weights between visual and textual tokens in order to spatially disentangle the customized concepts. We define a loss objective to measure the extent to which the customized concepts are spatially disentangled. As illustrated in Fig. 2, the noised latent map is updated with the objective in the first stage. In the second stage, the diffusion models take the updated noised latent map as input and is produced via semantic merging for the next denoising step.
We attempt to measure the extent to which the customized concepts are spatially disentangled with the help of the cross-attention maps generated at each step of the diffusion process. They can be regarded as features showing how strongly each latent image patch is associated with each text token embedding. Each cross-attention layer in the unet takes latent image features and text features as input. After linear layer , , , the attention maps are then:
@@
To spatially disentangle different customized concepts, we define another loss term to reduce overlap of the regions activated by different customized modules. For sub-prompt , we first aggregate the attention maps of the trigger tokens to get one attention map. Here, we empirically average the attention maps. After being smoothed by a Gaussian filter, we get the final attention map indicating the activated region of the customized concept. Figure 3 visualizes the attention maps of the involved concepts. Our proposed inter-prompt disentanglement loss is then:
The overall loss objective is a linear combination of and with coefficient , and is updated to with learning rate :
@@
Our method can be extended to enhance the compositional capabilities of text-to-image generation. The customized modules are omitted as customization is not needed in such setting. Besides and , we also include proposed by [8]. attempts to maximize the attention values for each subject token so as to encourage the occurrence of the subject. For a subject name that has multiple tokens, we empirically average the attention maps corresponding to each of them. Then the loss objective for compositional generation is defined as:
@@
3.4 Attention Grounding Training
Not all concepts have a clear correspondence between the trigger token and the activated area. As shown in the first row of Fig. 4, when the corresponding attention maps of <cat> and <chair> are somewhat messy, the MCG will not take the optimal effect. We incorporate two loss terms to explicitly strengthen the correspondence between the trigger tokens and the subject. The two loss terms are proposed by [52] to finetune the diffusion model for better performance on generating multiple subjects. We adapt them for single-concept customization learning.
In Eq. 12, represents a binary mask for the cross-attention map of the -th token. Equation 12 and Eq. 13 both encourage the model to have higher attention score within the mask region of the subjects. To enable the two loss terms, we produce masks for the reference images with SAM [25, 43]. The single-concept models are trained with a linear combination of the diffusion loss term, and . As shown in the second row of Fig. 4, the concepts have more clear attention maps.
@@ 4 Experiments
4.1 MC++ Benchmark
We construct a new benchmark for multi-concept customization, termed MC++. The previous benchmark CustomConcept101 [26] contains of 101 concepts with 3-15 images for each concept. For evaluating multi-concept customization, the CustomConcept101 contains prompts for 101 compositions of two subjects. Each composition has 12 corresponding prompts, e.g. photo of the {0} and {1}. However, the CustomConcept101 is not sufficient for evaluating compositions of more than two concepts. To evaluate the performance of customization methods on composition of more than two subjects, we collect prompts for compositions of three and four concepts. The MC++ contains 2,064 prompts in total. For details of the MC++, please refer to suppl.
@@ 4.2 Experiment Setup
Datasets. For customized multi-concept generation, we perform experiments on the MC++ benchmark.
@@
Evaluation metrics. For customized multi-concept generation, subject fidelity and prompt fidelity are evaluated, following [45, 26]. For subject fidelity, we compute two metrics: CLIP-I [14] and DINO [45]. CLIP-I is the average cosine similarity between CLIP [40] embeddings of generated and real images. The DINO metric is the average pairwise cosine similarity between the ViT-S/16 DINO [6] embeddings of generated and real images. For a composition of two concepts, CLIP-I and DINO metrics are computed for each concept respectively then averaged. Prompt fidelity is measured as the average cosine similarity between prompt and image CLIP embedding, denoted as CLIP-T[45]. The placeholder token in the text prompt template is replaced with the proper concept category.
@@
Baselines. For customized multi-concept generation, we compare our method with Custom Diffusion [26], Mix-of-Show [16], Cones 2 [31] and Textual Inversion [14]. Custom Diffusion optimizes the cross-attention key and value parameter in the unet plus a token embedding given reference images. Mix-of-Show trains ED-LoRA for each concept, then fuses multiple ED-LoRAs into one via gradient fusion. Mix-of-Show introduces regionally controllable sampling to control the location of each concept, which requires user-input bounding boxes for each concept, thus is omitted in our experiment. Cones 2 finetunes the text encoder for each concept. Then a residual token embedding is derived from the text encoder for inference. It proposes a layout guidance which requires user-input location information. The layout guidance is omitted in our experiment.
@@
For compositional generation, we compare our method with Attend-and-Excite [8], Divide-and-Bind [28] and Composable Diffusion [29]. Attend-and-Excite proposes a loss objective to maximize the attention values for each subject token. Divide-and-Bind proposes a attendance loss and a binding loss to better incorporate the semantics.
@@
Implementation details. We use the Stable Diffusion v1-5 in our experiments. Without loss of generality, we choose LoRA as the single-concept customized model, as different single-concept customization methods basically finetune some selected part of the diffusion model with similar loss functions. We use DPM++ sampler [32] with 30 sampling steps. Given a prompt template, e.g. photo of the {0} and {1}, we construct the sub-prompts for a composition as follows: (i) photo of the <cate1> and <cate2>, (ii) photo of the <concept1> and <cate2>, a <concept1>, (iii) photo of the <cate1> and <concept2>, a <concept2>. <cate1> is the category name of <concept1> defined by CustomConcept101, e.g. dog for pet_dog1. The detailed hyperparameter setting is in suppl.
Method | Four concepts | Three concepts | Two concepts | Overall | ||||||||
CLIP-T↑ | CLIP-I↑ | DINO↑ | CLIP-T↑ | CLIP-I↑ | DINO↑ | CLIP-T↑ | CLIP-I↑ | DINO↑ | CLIP-T↑ | CLIP-I↑ | DINO↑ | |
Custom Diffusion [26] | 0.753 | 0.602 | 0.236 | 0.800 | 0.606 | 0.269 | 0.789 | 0.645 | 0.339 | 0.781 | 0.618 | 0.281 |
Mix-of-Show [16] | 0.724 | 0.635 | 0.288 | 0.740 | 0.635 | 0.320 | 0.728 | 0.670 | 0.394 | 0.731 | 0.647 | 0.334 |
Cones 2 [31] | 0.807 | 0.618 | 0.244 | 0.806 | 0.602 | 0.244 | 0.792 | 0.629 | 0.295 | 0.802 | 0.616 | 0.258 |
Textual Inversion [14] | 0.656 | 0.618 | 0.244 | 0.753 | 0.621 | 0.280 | 0.746 | 0.648 | 0.322 | 0.718 | 0.629 | 0.282 |
Ours | 0.825 | 0.638 | 0.268 | 0.805 | 0.645 | 0.324 | 0.780 | 0.692 | 0.416 | 0.803 | 0.658 | 0.336 |
Method | Full Prompt↑ | Min. Object↑ | Text-Text↑ |